Determine the amounts of 20% and 50% salt solutions that should be mixed to obtain 300 gallons of 41% salt solution. Show all work.
Answer by Alan3354(69443)

Determine the amounts of 20% and 50% salt solutions that should be mixed to obtain 300 gallons of 41% salt solution.
------------
t = 20%
f = 50%
-----
t + f = 300 (total solution)
20t + 50f = 41*300 (total salt)
----
Solve for t & f